### v1

@@ 4.2 The Analysis
By item 1 of 4, $v$ is in at least one core of $\mathcal{R}$. Let $R\in\mathcal{R}$ be the core of the smallest rank containing $v$. Let $B$ be the center bag of $R$. Note that $Q(v)=B$ by definition.
@@
    All vertices in $U$ are equivalent under the tree ordering $\preceq$. This is because $Q(u)=B$ for every $u\in U$ by 3 and definition of $Q(\cdot)$.
@@
    For any two vertices $z\in R\setminus U$ and $u\in U$, $z\prec u$. To see this, observe that, by item 4 in 4, $z$ is in the attachment of some bag $Y$, which is the descendant of the center bag $B$. By item 2 and the definition of $Q(\cdot)$, $Q(z)$ is a (proper) descendant of $Y$. Therefore, $Q(z)$ is a descendant of $B$, which is $Q(u)$, implying that $z\prec u$.
@@
Suppose for contradiction that there exists $v\in R_{1}\cap R_{2}$. By item 3 in 4, $R_{k}$ only contains vertices in descendant bags of $B_{k}$. As $R_{1}\cap R_{2}\not=\emptyset$, either $B_{1}$ is an ancestor of $B_{2}$ or $B_{2}$ is ancestor of $B_{1}$. W.l.o.g., we assume that $B_{1}$ is an ancestor of $B_{2}$. Let $Y$ be the bag containing $v$. Then $Y$ is the descendant of both $B_{1}$ and $B_{2}$. As endpoints of edges of $G$ are either in the same bag or in two adjacent bags of the tree partition, $R_{1}\cap Y\not=\emptyset$ implies that $R_{1}\cap B_{2}\not=\emptyset$. Thus, $B_{2}$ will be marked as visited in algorithm 2 of the algorithm, and hence $R_{2}$ will not have the same rank as $R_{1}$, a contradiction. ∎
@@
Consider an arbitrary iteration where $B^{\prime}$ remains uncovered; that is, at least some vertex of $B^{\prime}$ is uncovered. Let $\mathcal{T}^{\prime}$ be the connected component of uncovered bags of $\mathcal{T}$ containing $B^{\prime}$. If $B^{\prime}$ got picked in algorithm 2, then all uncovered vertices of $B^{\prime}$ are marked as covered (in algorithm 2), and hence the claim holds. Otherwise, $B^{\prime}$ is marked visited in algorithm 2 when a core $R\in\mathcal{R}$ is created. Let $H$ be the support graph of $R$. By item 2 in 4, the attachment of every bag in $\mathcal{T}^{\prime}$ does not contain any vertex of $B^{\prime}$. Thus, $H$ only contains uncovered vertices of $B^{\prime}$. As $R\cap B^{\prime}\not=\emptyset$, $R$ contains at least one uncovered vertex of $B^{\prime}$, which will be marked as covered in algorithm 2; the claim holds. ∎
@@
  • Let $r_{1}$ be the lowest rank among all cores in $\mathfrak{C}$. We will show in 5 below that there is only one core in $\mathfrak{C}$ having rank $r_{1}$. Let this unique core in $\mathfrak{C}$ with rank $r_{1}$ be $R^{*}_{1}$. Let $\mathfrak{C}_{1}$ be the set of all cores in $\mathfrak{C}$ that are ancestors of $R^{*}_{1}$, including $R^{*}_{1}$. We then show in the next 6 that each core in $\mathfrak{C}_{1}$ intersects the center of $R^{*}_{1}$. Since the center is contained in a bag, Corollary 1 implies that $|\mathfrak{C}_{1}|\leq{\rm tp}^{2}$.
  • Next, let ¯$\bar{\mathfrak{C}}_{1}=\mathfrak{C}\setminus\mathfrak{C}_{1}$. If ¯$\bar{\mathfrak{C}}_{1}$ is non-empty, we define $r_{2}$ be the lowest rank among cores in $\bar{\mathfrak{C}_{1}}$. Note that $r_{2}>r_{1}$. 5 below again implies that there is only one core in ¯$\bar{\mathfrak{C}}_{1}$ having rank $r_{2}$. Let this unique core in ¯$\bar{\mathfrak{C}}_{1}$ with rank $r_{2}$ be $R^{*}_{2}$. Let $\mathfrak{C}_{2}$ be the set of all ancestor cores $R^{*}_{2}$ in ¯$\bar{\mathfrak{C}}_{1}$ including $R^{*}_{2}$. Then 6 implies that each core in $\mathfrak{C}_{2}$ intersects the center of $R^{*}_{2}$. Since the center is contained in a bag, Corollary 1 gives that $|\mathfrak{C}_{2}|\leq{\rm tp}^{2}$.
  • Inductively, we define the sequence of sets of cores $\mathfrak{C}_{1},\mathfrak{C}_{2},\mathfrak{C}_{3},\dots,\mathfrak{C}_{\ell}$ and $r_{1}<r_{2}<\dots<r_{\ell}$ until ¯=¯1subscript¯subscript¯$\bar{\mathfrak{C}}_{\ell}=\bar{\mathfrak{C}}_{\ell-1}\setminus\mathfrak{C}_{\ell}$ is empty. Here, $r_{j}$ is defined as the lowest rank among cores in ¯$\bar{\mathfrak{C}}_{j-1}$, and ¯$\bar{\mathfrak{C}}_{0}=\mathfrak{C}$. For each $j\in[\ell]$, 5 implies that there exist a unique core $R^{*}_{j}$ in ¯$\bar{\mathfrak{C}}_{j-1}$ with rank $r_{j}$. Let $\mathfrak{C}_{j}$ be the set of all cores in ¯$\bar{\mathfrak{C}}_{j}$ that are ancestors of $R^{*}_{j}$, including $R^{*}_{j}$. 6 then implies that all cores in $\mathfrak{C}_{j}$ intersect the center of $R^{*}_{j}$ in. Thus, $|\mathfrak{C}_{j}|\leq{\rm tp}^{2}$ for each $j\in[\ell]$ by Corollary 1. By Lemma 11, $r_{1}<r_{2}<\dots<r_{\ell}\leq{\rm tp}$, implying that $|\mathfrak{C}|\leq{\rm tp}^{3}$ as desired.
@@
Figure 5: Illustration for the proof of 5.
@@
The discussion in the proof outline with 5 and 6 proves the lemma. ∎
@@ 5 Applications
In this section, we gave a more detailed exposition of the applications of Theorem 1 and Theorem 2 mentioned in Section 1. The list of applications here is not meant to be exhaustive, and we believe that our result will find further applications.
@@ 5.8 Stochastic decomposition for minor-free graphs - a reduction from additive stretch embeddings
$\Pr\left[\Psi_{1}\right]=\Pr\left[\mathbf{B}_{G}(v,\frac{\Delta}{t})\subseteq C\right]=\Pr\left[\mathbf{B}_{G}(v,\frac{\Delta}{t\Delta^{\prime}}\cdot\Delta^{\prime})\subseteq C\right]\geq e^{-O(r)\cdot\frac{\Delta}{t\Delta^{\prime}}}\geq\frac{4}{5}\leavevmode\nobreak\ .$
@@
$\leq\frac{{\mathbb{E}}\left[d_{H_{C}}(u,v)-d_{G[C]}(u,v)\right]}{5\cdot\epsilon\cdot\Delta^{\prime}}\leq\frac{1}{5}\leavevmode\nobreak\ .$
@@
$\Pr\left[\Psi_{3}\mid\Psi_{1}\wedge\Psi_{2}\right]\geq\Pr\left[\mathbf{B}_{H_{C}}(v,\frac{2\Delta}{t})\subseteq\mathcal{P}_{C}(c)\right]\geq e^{-O(\log\pi(\Omega(\frac{1}{r}),r,n))\cdot\frac{2}{t}}\geq\frac{4}{5}\leavevmode\nobreak\ ,$
@@
$\Pr\left[\mathcal{P}(v)=\mathcal{P}(u)\right]\geq\Pr\left[\Psi_{1}\wedge\Psi_{2}\wedge\Psi_{3}\right]\geq\left(\frac{4}{5}\right)^{3}>\frac{1}{2}\leavevmode\nobreak\ .$

### v2

@@ 4.2 The Analysis
By item 1 of 20, $v$ is in at least one core of $\mathcal{R}$. Let $R\in\mathcal{R}$ be the core of the smallest rank containing $v$. Let $B$ be the center bag of $R$. Note that $Q(v)=B$ by definition.
@@
    All vertices in $U$ are equivalent under the tree ordering $\preceq$. This is because $Q(u)=B$ for every $u\in U$ by 22 and definition of $Q(\cdot)$.
@@
    For any two vertices $z\in R\setminus U$ and $u\in U$, $z\prec u$. To see this, observe that, by item 4 in 20, $z$ is in the attachment of some bag $Y$, which is the descendant of the center bag $B$. By item 2 and the definition of $Q(\cdot)$, $Q(z)$ is a (proper) descendant of $Y$. Therefore, $Q(z)$ is a descendant of $B$, which is $Q(u)$, implying that $z\prec u$.
@@
Suppose for contradiction that there exists $v\in R_{1}\cap R_{2}$. By item 3 in 20, $R_{k}$ only contains vertices in descendant bags of $B_{k}$. As $R_{1}\cap R_{2}\not=\emptyset$, either $B_{1}$ is an ancestor of $B_{2}$ or $B_{2}$ is ancestor of $B_{1}$. W.l.o.g., we assume that $B_{1}$ is an ancestor of $B_{2}$. Let $Y$ be the bag containing $v$. Then $Y$ is the descendant of both $B_{1}$ and $B_{2}$. As endpoints of edges of $G$ are either in the same bag or in two adjacent bags of the tree partition, $R_{1}\cap Y\not=\emptyset$ implies that $R_{1}\cap B_{2}\not=\emptyset$. Thus, $B_{2}$ will be marked as visited in algorithm 2 of the algorithm, and hence $R_{2}$ will not have the same rank as $R_{1}$, a contradiction. ∎
@@
Consider an arbitrary iteration where $B^{\prime}$ remains uncovered; that is, at least some vertex of $B^{\prime}$ is uncovered. Let $\mathcal{T}^{\prime}$ be the connected component of uncovered bags of $\mathcal{T}$ containing $B^{\prime}$. If $B^{\prime}$ got picked in algorithm 2, then all uncovered vertices of $B^{\prime}$ are marked as covered (in algorithm 2), and hence the claim holds. Otherwise, $B^{\prime}$ is marked visited in algorithm 2 when a core $R\in\mathcal{R}$ is created. Let $H$ be the support graph of $R$. By item 2 in 20, the attachment of every bag in $\mathcal{T}^{\prime}$ does not contain any vertex of $B^{\prime}$. Thus, $H$ only contains uncovered vertices of $B^{\prime}$. As $R\cap B^{\prime}\not=\emptyset$, $R$ contains at least one uncovered vertex of $B^{\prime}$, which will be marked as covered in algorithm 2; the claim holds. ∎
@@
  • Let $r_{1}$ be the lowest rank among all cores in $\mathfrak{C}$. We will show in 31 below that there is only one core in $\mathfrak{C}$ having rank $r_{1}$. Let this unique core in $\mathfrak{C}$ with rank $r_{1}$ be $R^{*}_{1}$. Let $\mathfrak{C}_{1}$ be the set of all cores in $\mathfrak{C}$ that are ancestors of $R^{*}_{1}$, including $R^{*}_{1}$. We then show in the next 32 that each core in $\mathfrak{C}_{1}$ intersects the center of $R^{*}_{1}$. Since the center is contained in a bag, Corollary 27 implies that $|\mathfrak{C}_{1}|\leq{\rm tp}^{2}$.
  • Next, let ¯$\bar{\mathfrak{C}}_{1}=\mathfrak{C}\setminus\mathfrak{C}_{1}$. If ¯$\bar{\mathfrak{C}}_{1}$ is non-empty, we define $r_{2}$ be the lowest rank among cores in $\bar{\mathfrak{C}_{1}}$. Note that $r_{2}>r_{1}$. 31 below again implies that there is only one core in ¯$\bar{\mathfrak{C}}_{1}$ having rank $r_{2}$. Let this unique core in ¯$\bar{\mathfrak{C}}_{1}$ with rank $r_{2}$ be $R^{*}_{2}$. Let $\mathfrak{C}_{2}$ be the set of all ancestor cores $R^{*}_{2}$ in ¯$\bar{\mathfrak{C}}_{1}$ including $R^{*}_{2}$. Then 32 implies that each core in $\mathfrak{C}_{2}$ intersects the center of $R^{*}_{2}$. Since the center is contained in a bag, Corollary 27 gives that $|\mathfrak{C}_{2}|\leq{\rm tp}^{2}$.
  • Inductively, we define the sequence of sets of cores $\mathfrak{C}_{1},\mathfrak{C}_{2},\mathfrak{C}_{3},\dots,\mathfrak{C}_{\ell}$ and $r_{1}<r_{2}<\dots<r_{\ell}$ until ¯=¯1subscript¯subscript¯$\bar{\mathfrak{C}}_{\ell}=\bar{\mathfrak{C}}_{\ell-1}\setminus\mathfrak{C}_{\ell}$ is empty. Here, $r_{j}$ is defined as the lowest rank among cores in ¯$\bar{\mathfrak{C}}_{j-1}$, and ¯$\bar{\mathfrak{C}}_{0}=\mathfrak{C}$. For each $j\in[\ell]$, 31 implies that there exist a unique core $R^{*}_{j}$ in ¯$\bar{\mathfrak{C}}_{j-1}$ with rank $r_{j}$. Let $\mathfrak{C}_{j}$ be the set of all cores in ¯$\bar{\mathfrak{C}}_{j}$ that are ancestors of $R^{*}_{j}$, including $R^{*}_{j}$. 32 then implies that all cores in $\mathfrak{C}_{j}$ intersect the center of $R^{*}_{j}$ in. Thus, $|\mathfrak{C}_{j}|\leq{\rm tp}^{2}$ for each $j\in[\ell]$ by Corollary 27. By Lemma 26, $r_{1}<r_{2}<\dots<r_{\ell}\leq{\rm tp}$, implying that $|\mathfrak{C}|\leq{\rm tp}^{3}$ as desired.
@@
Figure 5: Illustration for the proof of 31.
@@
The discussion in the proof outline with 31 and 32 proves the lemma. ∎
@@ 5 Applications
In this section, we give a more detailed exposition of the applications of Theorem 2 and Theorem 3 mentioned in Section 1. The list of applications here is not meant to be exhaustive, and we believe that our result will find further applications.
@@ 5.8 Stochastic decomposition for minor-free graphs - a reduction from additive stretch embeddings
$\Pr\left[\Psi_{1}\right]=\Pr\left[\mathbf{B}_{G}(v,\frac{\Delta}{t})\subseteq C\right]=\Pr\left[\mathbf{B}_{G}(v,\frac{\Delta}{t\Delta^{\prime}}\cdot\Delta^{\prime})\subseteq C\right]\geq e^{-O(r)\cdot\frac{\Delta}{t\Delta^{\prime}}}\geq\frac{4}{5}~{}.$
@@
$\leq\frac{{\mathbb{E}}\left[d_{H_{C}}(u,v)-d_{G[C]}(u,v)\right]}{5\cdot\epsilon\cdot\Delta^{\prime}}\leq\frac{1}{5}~{}.$
@@
$\Pr\left[\Psi_{3}\mid\Psi_{1}\wedge\Psi_{2}\right]\geq\Pr\left[\mathbf{B}_{H_{C}}(v,\frac{2\Delta}{t})\subseteq\mathcal{P}_{C}(c)\right]\geq e^{-O(\log\pi(\Omega(\frac{1}{r}),r,n))\cdot\frac{2}{t}}\geq\frac{4}{5}~{},$
@@
$\Pr\left[\mathcal{P}(v)=\mathcal{P}(u)\right]\geq\Pr\left[\Psi_{1}\wedge\Psi_{2}\wedge\Psi_{3}\right]\geq\left(\frac{4}{5}\right)^{3}>\frac{1}{2}~{}.$
Acknowledgements.
Arnold Filtser was supported by the Israel Science Foundation (grant No. 1042/22). Hung Le was supported by the NSF CAREER Award No. CCF-223728, an NSF Grant No. CCF-2121952, and a Google Research Scholar Award.